Consider a linear regression model where y represents the response variable, x is a quantitative explanatory variable, and d is a dummy variable. The model is estimated as  yˆy^  = 14.6 + 4.5x − 3.4d.

a. Interpret the dummy variable coefficient.

  • Intercept shifts down by 3.4 units as d changes from 0 to 1.

  • Slope shifts down by 3.4 units as d changes from 0 to 1.

  • Intercept shifts up by 3.4 units as d changes from 0 to 1.

  • Slope shifts up by 3.4 units as d changes from 0 to 1.

b. Compute y^ for x = 2 and d = 1. (Round your answer to 1 decimal place.)

c. Compute y^  for x = 2 and d = 0. (Round your answer to 1 decimal place.)
Answer #1

a)Intercept shifts down by 3.4 units as d changes from 0 to 1

b) from given regression equation: y^ =14.6+4.5*2-3.4*1 =20.2

c) y^ =14.6+4.5*2-3.4*0 =23.6